OpenStudy (anonymous):

PLEASE HELP I WILL GIVE MEDAL Divide 3x-15/14y / x^2-15x+50/7

OpenStudy (anonymous):

is it two fractions being divided?

OpenStudy (anonymous):

it is \[\frac{ 3x - 15}{ 14y} \div \frac{ x ^{2} -15x+50}{ 7 }\]

OpenStudy (anonymous):

\[\frac{ 3x-15 }{ 14y }\div \frac{ x ^{2}-15x+50 }{ 7}\]

OpenStudy (anonymous):

ok just a second (:

OpenStudy (anonymous):

ok thank you so much!

OpenStudy (anonymous):

first you have to flip the second fraction, making it 3x−15 /14y times 7/x 2 −15x+50

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now you factor, so you should get \[\frac{ 3(x-5)}{ 14y} \times \frac{ 7 }{(x-5)(x-10)} \]

OpenStudy (anonymous):

now you have to cancel out the things you can, (x-5) and (x-5), and 7 and 14. now you have \[\frac{ 3 }{ 2y } \times \frac{ 1 }{ (x-10) }\]

OpenStudy (anonymous):

okat do you cross multiply

OpenStudy (anonymous):

nope, the answer is 3 / 2y(x-10) im pretty sure. is that one of the choices??

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

can you help me with one more

OpenStudy (anonymous):

\[\frac{ \left( \ \frac{ 18v }{ t ^{2} - 4t + 3} \right) }{ \frac{ \left( 3v ^{2}\right)}{ 6t- 6} }\]

OpenStudy (anonymous):

@indigochild